OpenStudy (anonymous):

Please help i have so much more work to do today. What is the vertex of the graph of y = -1(x - 2)2 + 3? i got (-2, 3) is that correct?

OpenStudy (jdoe0001):

is already in vertex form, so the vertex is at \(\bf y = -1(x - \color{red}{2})^2 \color{red}{+ 3}\) notice that \(\bf y = -1(x - 2)^2 + 3\implies y = -1(x - (+2))^2 + 3\)

OpenStudy (jdoe0001):

for example, a vertex of ( -2, 3) will be \(\bf y = -1(x + 2)^2 + 3\implies y = -1(x - (-2))^2 + 3\)

OpenStudy (anonymous):

@jdoe0001 i think i get it.....well kinda it only half clicks

OpenStudy (anonymous):

i dont understand what add the extra )) did?

OpenStudy (anonymous):

sorry i feel dumb :(

OpenStudy (anonymous):

@jdoe0001 So if i under stand right, i got it right?

OpenStudy (jdoe0001):

ahemm.... well... if we 're referring to \(\bf y = -1(x + 2)^2 + 3\implies y = -1(x - (-2))^2 + 3\) yes you did if we're referring to \(\bf y = -1(x - \color{red}{2})^2 \color{red}{+ 3}\) you're pretty close

OpenStudy (anonymous):

Oh im dumb thanks lol

OpenStudy (jdoe0001):

the vertex for is \(\bf y = a(x-h)^2+k\) where "h" is the x-coordinate, and the "k" is the y-coordinate so based on the template, "h" is just +2, not minus, because the minus is just part of the vertex form, not part of the value itself